OpenStudy (anonymous):

Can someone show me the steps to solving this equation? 5(d+4)=7(d-2)

OpenStudy (anonymous):

5*d+5*4 =7*d-7*2 5d+20 = 7d-14 2d =34 d =17

OpenStudy (anonymous):

5d+20 =7d-14 34 = 2d 17 = d

OpenStudy (anonymous):

You lost be at 2d=34? how did you get 2d?

OpenStudy (anonymous):

me*

OpenStudy (anonymous):

well, I subtracted 5d from both sides, and added 14 to both sides

OpenStudy (anonymous):

Oh, ok